OpenStudy (anonymous):

so lost, could someone work this out for me: function y=cos x/5 on the interval [0,10pi] specify the int. in which the function is increasing.

OpenStudy (anonymous):

is it y=(cosx)/5 or cos(x/5)

OpenStudy (anonymous):

cos (x/5)

OpenStudy (anonymous):

Strictly increasing:\[5\pi <x<10\pi \]

OpenStudy (anonymous):

robto. how did u get that tho?